# L4Q: Parameter Efficient Quantization-Aware Fine-Tuning on Large Language Models

### v5

@@ Abstract
Hyesung Jeon Seoul National University hjeon2k@snu.ac.kr                        Yulhwa Kim Sungkyunkwan University yulhwakim@skku.edu                        Jae-Joon Kim Seoul National University kimjaejoon@snu.ac.kr
@@ 1 Introduction
Concurrently, to enhance the problem-solving abilities of LLMs for specific applications, fine-tuning pre-trained LLMs on downstream tasks is crucial as it improves accuracy on target tasks and related tasks (Wei et al., 2022; Scialom et al., 2022). However, fine-tuning is a resource-intensive process due to the large number of model weight parameters involved. Parameter-efficient fine-tuning (PEFT) addresses this issue (Hu et al., 2022; Li and Liang, 2021; Liu et al., 2022a; Wang et al., 2023) by training a small subset of parameters while freezing the majority of pre-trained weights. One of the most prominent techniques within PEFT is Low-Rank Adaptation (LoRA) (Hu et al., 2022), which inserts trainable rank decomposition matrices into each layer to represent updates to the frozen weights.
@@ 3.1 Straight Integration of QAT and LoRA
Limitations of QAT-LoRA  However, several issues arise with this straightforward integration of QAT and LoRA, where quantized weights and LoRA parameters remain distinct. First, although freezing the pre-trained weights eliminates the need for optimizer states, weight gradients $\frac{\partial L}{\partial W_{q}}$ must still be stored to update the quantization parameters, as shown in Equation 5 and Equation 6. As a result, QAT-LoRA still incurs memory overhead from weight gradients, undermining the memory efficiency benefits of LoRA fine-tuning. Secondly, QAT-LoRA produces a mixed-precision model with both quantized weights and high-precision LoRA parameters. This mixed-precision approach negates the advantages of LLM quantization, similar to previous methods such as QLoRA and LoftQ discussed in Section 2.4. Lastly, the gradient updates for quantization and LoRA parameters are decoupled, limiting the potential for comprehensive optimization across the model. As outlined in Equations  5 and  6, updates to the quantization parameters rely on the quantized weight matrix $W_{q}$, while updates to the LoRA parameters depend on weights $A$ and $B$. This limits the effectiveness of model training, as it prevents holistic adjustments where changes in LoRA parameters could directly influence quantization adjustments and vice versa.
@@ 4.1 Experimental Settings
Baselines  We compare the proposed L4Q with previous quantization methods and quantization-aware PEFT methods. The baseline quantization methods considered are LSQ for QAT, and GPTQ and OmniQuant for PTQ. For quantization-aware PEFT baselines, we include QLoRA, QA-LoRA, and LoftQ. We apply uniform quantization in our experiments to ensure consistency across methods. The methods that were originally deployed with non-uniform quantization are denoted with an asterisk (*).
Evaluation Setups  We establish the L4Q framework based on the open-source frameworks: Lit-GPT [^2] https://github.com/Lightning-AI/lit-gpt.git and huggingface transformers [^3] https://github.com/huggingface/transformers.git. The models are symmetrically quantized with quantization group size of 128 for the LLaMA and Mistral models, and 64 for the OpenLLaMA models due to its small channel size. Also, the models are fine-tuned with a LoRA rank $r$ of 4 by default, and 8 for Mistral. We double the rank $r$ in QA-LoRA for a fair comparison in terms of the number of LoRA parameters. Further details on the effect of rank size and quantization group size can be found in Appendix C. We use the Stanford-Alpaca (Taori et al., 2023), an English based dataset that consists of 50k training samples and 2k validation samples generated from the GPT 3.5 model (Brown et al., 2020). We use bfloat16 precision for numerically stable fine-tuning. All experiments are conducted on an NVIDIA A100 80GB GPU. Detailed hyperparameter settings for fine-tuning are provided in Appendix D.
Evaluation Metrics  We evaluate the accuracy of LLMs on Commonsense QA (CSQA) (Gao et al., 2021) and MMLU (Hendrycks et al., 2021) benchmarks. The CSQA benchmark comprises seven representative multiple-choice tasks designed to evaluate language models (Zellers et al., 2019; Bisk et al., 2020; Clark et al., 2018; Sakaguchi et al., 2020; Clark et al., 2019; Talmor et al., 2019). The MMLU benchmark spans four subject categories made up of 57 subcategories of language tasks.
@@ 4.2 Evaluation Results
Inference Speedup  We measure the inference speed of 16-bit pre-trained models and quantized models using LLaMA-1 models. The quantized models include fully-quantized 4-bit models (L4Q and QA-LoRA), which contain only quantized parameters, and mixed-precision 4&16-bit models (LoftQ*, QLoRA*, and QAT-LoRA), which use additional 16-bit LoRA parameters. The inference speed was measured with input batch sizes ranging from 1 to 64. The average speedup of quantized models compared to full-precision 16-bit models is reported in Figure 4. The 4-bit models achieve a speedup of 1.8$\times$ to 2.3$\times$ over the 16-bit models. More importantly, these 4-bit models achieve a 1.4$\times$ to 1.6$\times$ speedup compared to mixed-precision 4&16-bit models, which are also quantized versions of LLMs. This demonstrates that the full integration of QAT and LoRA in L4Q plays a crucial role in achieving the best inference speed. More analysis on speedup can be found in Appendix F.
@@ 5 Conclusion
Our work focuses on efficient weight quantization methods for large language models (LLMs), but there are additional avenues that could further enhance inference efficiency and effectiveness, particularly when combined with L4Q. Activation quantization offers a promising way to further reduce memory and computation costs when combined with weight quantization, potentially optimizing LLM performance. Similarly, KV cache compression through pruning or quantization could minimize memory overhead and latency, especially for long-context applications. Finally, refinement of LoRA initialization schemes for quantized models may enhance fine-tuning efficiency and improve performance in low-resource settings.
Although we have not investigated these vertical methods within the context of L4Q, we believe that these directions offer promising opportunities for future research. Integrating these approaches with L4Q could further improve LLM inference efficiency and effectiveness, which we leave to future work.
@@
This paper presents work aimed at advancing machine learning, particularly in the context of language models. While our research contributes to the development and application of LLMs, we recognize the potential societal implications associated with this work. Based on the claims of the referenced sources — including datasets, code, and models — we believe that the artifacts used and the models produced do not violate personal identification rights or contain offensive content. Consequently, we do not further examine or discuss potential harms or biases inherent in these models.
All datasets, code, and models cited in this paper are publicly accessible and processed solely for research purposes. Our use of these artifacts is consistent with their intended purpose.
@@ Appendix G Detailed Result on Main Evaluations
Model Method #Bits Hella. PIQA ARC-c ARC-e Winogr. BoolQ OBQA Avg.

### v6

@@ Abstract
Hyesung Jeon1  Yulhwa Kim2thanks: Corresponding author.  Jae-Joon Kim1∗ 1Seoul National University  2Sungkyunkwan University {hjeon2k, kimjaejoon}@snu.ac.kryulhwakim@skku.edu
@@ 1 Introduction
Concurrently, to enhance the problem-solving abilities of LLMs for specific applications, fine-tuning pre-trained LLMs on downstream tasks is crucial as it improves accuracy on target tasks and related tasks (Wei et al., 2022; Scialom et al., 2022). However, fine-tuning is a resource-intensive process due to the large number of model weights involved. Parameter-efficient fine-tuning (PEFT) addresses this issue (Hu et al., 2022; Li and Liang, 2021; Liu et al., 2022a; Wang et al., 2023) by training a small subset of parameters while freezing the majority of pre-trained weights. One of the most prominent techniques within PEFT is Low-Rank Adaptation (LoRA) (Hu et al., 2022), which inserts trainable rank decomposition matrices into each layer to represent updates to the frozen weights.
@@ 3.1 Straight Integration of QAT and LoRA
Limitations of QAT-LoRA  While the previous section introduced a straightforward integration of QAT and LoRA—where quantized weights and LoRA parameters are maintained separately—this approach presents several limitations. First, although freezing the pre-trained weights eliminates the need for optimizer states, weight gradients $\frac{\partial L}{\partial W_{q}}$ must still be stored to update the quantization parameters, as shown in Equation 5 and Equation 6. As a result, QAT-LoRA still incurs memory overhead from weight gradients, undermining the memory efficiency benefits of LoRA fine-tuning. Secondly, QAT-LoRA produces a mixed-precision model with both quantized weights and high-precision LoRA parameters. This mixed-precision approach negates the advantages of LLM quantization, similar to previous methods such as QLoRA and LoftQ discussed in Section 2.4. Lastly, the gradient updates for quantization and LoRA parameters are decoupled, limiting the potential for comprehensive optimization across the model. As outlined in Equations  5 and  6, updates to the quantization parameters rely on the quantized weight matrix $W_{q}$, while updates to the LoRA parameters depend on weights $A$ and $B$. This limits the effectiveness of model training, as it prevents holistic adjustments where changes in LoRA parameters could directly influence quantization adjustments and vice versa.
@@ 4.1 Experimental Settings
Baselines  We compare the proposed L4Q with previous quantization methods and quantization-aware PEFT methods. The baseline quantization methods considered are LSQ for QAT, and GPTQ and OmniQuant for PTQ. For quantization-aware PEFT baselines, we include QLoRA, QA-LoRA, and LoftQ. We apply uniform quantization in our experiments to ensure consistency across methods. The methods that were originally deployed with non-uniform quantization are denoted with an asterisk (*). For methods with LoRA adapters, including L4Q, the adapters are initialized following the original scheme: $A$ is initialized using Kaiming-uniform, while $B$ is initialized to zeros. We also note that the term fully-quantized refers to a condition in which all linear layers are quantized, while non-linear components and the head layer remain in higher precision.
Evaluation Setups  We establish the L4Q framework based on the open-source frameworks: Lit-GPT [^2] https://github.com/Lightning-AI/lit-gpt.git and huggingface transformers [^3] https://github.com/huggingface/transformers.git. The models are symmetrically quantized with quantization group size of 128 for the LLaMA and Mistral models, and 64 for the OpenLLaMA models due to its small channel size. Also, the models are fine-tuned with a LoRA rank $r$ of 4 by default, and 8 for Mistral. We double the rank $r$ in QA-LoRA for a fair comparison in terms of the number of LoRA parameters. Further details on the effect of rank size and quantization group size can be found in Appendix C. We use the Stanford-Alpaca (Taori et al., 2023), a dataset that consists of 50k training samples and 2k validation samples generated from the GPT 3.5 (Brown et al., 2020). We use bfloat16 precision for numerically stable fine-tuning. All experiments are conducted on an NVIDIA A100 80GB GPU. Detailed hyperparameter settings for fine-tuning are provided in Appendix D.
Evaluation Metrics  We evaluate the accuracy of LLMs on Commonsense QA (CSQA) (Gao et al., 2021) and MMLU (Hendrycks et al., 2021) benchmarks. The CSQA benchmark comprises seven multiple-choice tasks designed to evaluate language models (Zellers et al., 2019; Bisk et al., 2020; Clark et al., 2018; Sakaguchi et al., 2020; Clark et al., 2019; Talmor et al., 2019). The MMLU benchmark spans four subject categories made up of 57 subcategories of language tasks.
@@ 4.2 Evaluation Results
Inference Speedup  We measure the inference speed of 16-bit pre-trained models and quantized models using LLaMA-1 models. The average speedup of quantized models compared to full-precision 16-bit models is reported in Figure 4. The quantized models include fully-quantized 4-bit models (L4Q and QA-LoRA), which contain only quantized parameters, and mixed-precision 4&16-bit models (LoftQ*, QLoRA*, and QAT-LoRA), which use additional 16-bit LoRA parameters. The inference speed was measured with input batch sizes ranging from 1 to 64. The 4-bit models achieve a speedup of 1.8$\times$ to 2.3$\times$ over the pre-trained models. Moreover, these 4-bit models achieve a 1.4$\times$ to 1.6$\times$ speedup compared to mixed-precision models, which are also quantized versions of LLMs. This demonstrates that the full integration of QAT and LoRA in L4Q plays a crucial role in inference speedup. Further analysis and details on speedup can be found in Appendix F.
@@ 5 Conclusion
Our work focuses on efficient weight quantization methods for large language models (LLMs), but there are vertical approaches that could further enhance inference efficiency and effectiveness, integrated with our work. Activation quantization offers a chance to further reduce computation costs when combined with weight quantization. Similarly, KV cache compression could minimize memory overhead and latency, especially for long-context applications. Finally, refinement of LoRA initialization schemes for quantized models may improve accuracy of the fine-tuned models. We believe that integrating these approaches with L4Q could further improve LLM inference efficiency and effectiveness, which we leave to future work.
@@
While our research contributes to the development and application of machine learning, particularly in language models, we recognize the potential societal implications associated with this work. Based on the claims of the referenced sources — including datasets, code, and models — we believe that the artifacts do not violate personal identification rights or contain offensive content. All datasets, code, and models cited in this paper are publicly accessible and processed solely for research purposes. Our use of these artifacts is consistent with their intended purpose.
Acknowledgments
This work was supported in part by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (No.RS-2025-02273157: Development of Low Power Training/Inference Technologies based on AI Semiconductors, RS-2024-00395134, DPU-Centric Datacenter Architecture for Next-Generation AI Devices, No. 2021-0-01343: Artificial Intelligence Graduate School Program (Seoul National University)), Samsung Research Funding Center under Project SRFC-TC1603-53, and BK21 FOUR program at Seoul National University. (Corresponding Author: Jae-Joon Kim).
Appendix
@@ Appendix G Detailed Result on Main Evaluations
LLaMA-3 8B None 16 60.2 79.7 50.4 80.1 72.5 81.4 34.8 65.6
LoRA 16 60.6 79.9 53.8 82.7 74.7 82.8 36.0 67.2
GPTQ 4 56.8 69.3 31.8 62.6 73.1 78.5 33.0 57.9
OmniQ 4 59.2 78.6 49.6 79.7 72.5 81.1 33.4 64.9
LoftQ* 4&16 57.9 78.7 46.9 78.4 70.3 77.5 33.6 63.3
QLoRA* 4&16 56.7 69.5 33.3 66.3 72.9 78.4 33.0 58.6
QA-LoRA 4 56.3 69.3 32.3 65.7 72.3 78.6 31.6 58.0
QAT-LoRA 4&16 59.1 79.8 50.1 80.2 74.0 83.0 34.2 65.7
L4Q 4 60.5 80.4 52.7 81.6 73.6 83.6 35.0 66.8
GPTQ 3 51.8 68.8 30.2 58.3 67.7 70.1 27.4 53.5
OmniQ 3 55.0 76.7 39.2 69.2 69.2 72.6 28.8 58.7
LoftQ* 3&16 35.9 68.8 29.8 57.7 59.0 67.9 20.8 48.6
QLoRA* 3&16 53.4 72.2 31.8 62.9 69.2 71.5 28.8 55.7
QA-LoRA 3 52.7 73.5 36.4 67.9 67.6 71.7 26.6 56.6
QAT-LoRA 3&16 56.6 78.2 47.4 77.8 68.0 80.6 33.4 63.1
L4Q 3 56.5 78.1 47.8 78.6 69.2 82.0 32.2 63.5
@@
Model Method #Bits Hella. PIQA ARC-c ARC-e Winogr. BoolQ OBQA Avg.